OpenStudy (anonymous):

(X³-x/2x-6)(x²-2x-3/x²-1)

jhonyy9 (jhonyy9):

you have wrote it right sure ?

jhonyy9 (jhonyy9):

x2 -2x-3=(x-3)(x+1) x2 -1=(x-1)(x+1) x3 -x=x(x2 -1)=x(x-1)(x+1) 2x-6=2(x-3)

jhonyy9 (jhonyy9):

check what and how you will can simplify